OpenStudy (anonymous):

Eliminate the parameter x=4+2cos(theta) Y=1+2sin(theta)

OpenStudy (anonymous):

I dont understand thi problem

OpenStudy (anonymous):

Do i have to solve for x and then what?

OpenStudy (aravindg):

Separate sin theta and cos theta first.

OpenStudy (anonymous):

X-2=4cos(theta)

OpenStudy (anonymous):

X-2/4=cos(theta)

OpenStudy (aravindg):

It is x-4/2 mot x-2/4

OpenStudy (aravindg):

*not

OpenStudy (anonymous):

Am i rigtht

OpenStudy (anonymous):

No? :(

OpenStudy (aravindg):

(x-4)/2=cos theta

OpenStudy (anonymous):

So what can i do?

OpenStudy (anonymous):

Ok